Find a linear function h, given h(5)= -21 and h(-4)= 24. Then find h(2).
Answer by nerdybill(7384)

Find a linear function h, given h(5)= -21 and h(-4)= 24. Then find h(2).
.
"h(5)= -21 and h(-4)= 24" gives you two points:
(5, -21) and (-4,24)
Slope then is
m = (y2-y1)/(x2-x1)
m = (24-(-21))/(-4-5)
m = (24+21)/(-9)
m = (45)/(-9)
m = -5
.
plug slope (-5) and one point (5,-21) into the "point-slope" form:
y - y1 = m(x - x1)
y - (-21) = -5(x - 5)
y + 21 = -5x + 25
y = -5x + 4
.
h(2) then is:
h(2) = -5(2) + 4
h(2) = -10 + 4
h(2) = -6